OpenStudy (anonymous):

HELP ME! Enter the quadratic equation in standard form in the box below. x^2-5x-5=2x+3 (Please type in the actual answer, not how to do it.)

OpenStudy (shubhamsrg):

the actual ans is : "do your homework yourself" !! these ques are pathetic,,i understand you couldnt do 1st ques,,but afterwards!! amazing!!

OpenStudy (anonymous):

Dude seriously? Can somebody else "help" me? Thanks (:

OpenStudy (anonymous):

seriously you want the answer? Then pay me.

OpenStudy (anonymous):

Either that or take advantage of this free service and learn how to do it.

OpenStudy (shubhamsrg):

i wish no one helps you on this one.. you're indeed ,,helpless!!

jhonyy9 (jhonyy9):

x^2-5x-5=2x+3 add to both sides -2x-3 x^2-5x-5-2x-3=2x+3-2x-3 x^2-7x-8=0 this can being factorized (x+1)(x-8)=0 x+1=0 x=-1 x-8=0 x=8

Parth (parthkohli):

1) Subtract 2x from both sides. 2) Bring 0 on the right side by subtracting 3 from both sides.

OpenStudy (anonymous):

@jhonyy9 the asker is not asking you to solve it.

OpenStudy (anonymous):

Thanks jhonnyy9 & ParthKohli! (:

OpenStudy (anonymous):

Hahaha so true.

OpenStudy (shubhamsrg):

i know whats gonna be your next ques: solve for x : x=8 ..